OpenStudy (anonymous):

What is the third term of (a+b)^11

OpenStudy (cwrw238):

(r+1)th term = nCr a^(n-r) b^r

OpenStudy (cwrw238):

here r = 2, n = 11 3rd term = 11C2 a^9 b^2 - can you continue from here?

OpenStudy (anonymous):

11C2 a^9 b^2 is the answer?

OpenStudy (cwrw238):

well we need to calculate 11C2 - which is the number of combinations of 2 in 11 = 11* 10 ----- = 55 2 * 1 sao 3rd term is 55a^9 b^2

OpenStudy (anonymous):

Thank you for your help

OpenStudy (cwrw238):

yw